draw a rectangle that has an area of 6 square centimetres and a perimeter of 14 centimetres.
Answer by TX911(36)

A=LW
(x)(y)=6
Perimeter=2(w)+2(L)
2(w)+2(L)=16
This one you can see can be done in your head without much work:

6x1=6
and
6(2)+1(2)=14
So our length would be 6cm and our width would be 1cm
I can't draw on here but I'm sure you can take it from here.
Hope this helps